OpenStudy (anonymous):

Find the distance between parallel lines y=5x+2 and y=5x-7.

OpenStudy (akashdeepdeb):

First determine a convinient point on either line! Let us say we have the first one And the point can be (1,7) [BECAUSE IT FITS THE EQUATION] Okay so that being said we move on to our perpendicular! Now you should know that the slopes of Perpendicular lines are related as m1.m2 = -1 And Parrallel lines obviously always have the same slope!

OpenStudy (akashdeepdeb):

So the slope of the parrallel lines in these equations are -1/5 Okay so now we have a point (1,7) and slope of perpendicular as -1/5 Use equation = (y-y1) = m(x-x1) to find the perpendicular equation!!!

OpenStudy (akashdeepdeb):

After you are done with that use the perpendicular line and the other parrallel line to find their solution! [YOU GET THEIR INTERSECTION] After you get the two points where the perpendicular intersects with the 2 parrallel lines Use the distance formula [DISTANCE BETWEEN 2 POINTS] To get the shortest distance/Perpendicular between the 2 parralel lines! Understood?

OpenStudy (akashdeepdeb):

Refer to this diagram.

OpenStudy (akashdeepdeb):



OpenStudy (akashdeepdeb):

Understand? :D

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

would the distance be 4?

OpenStudy (anonymous):

Another way to go about this is by using the distance formula directly: \[d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] Here, \((x_1,y_1)\) is a point on one line (\(y=5x+2\)), and \((x_2,y_2)\) is a point on the other line (\(y=5x-7\)). So your first point is equivalent to \((x_1,5x_1+2)\) and your second point is \((x_2,5x_2-7)\): \[d=\sqrt{(x_1-x_2)^2+(5x_1+2-(5x_2-7))^2}\\ d=\sqrt{(x_1-x_2)^2+(5x_1-5x_2+9)^2}\] Then it's only a matter of determining the x-coordinates of the intersections of the parallel lines and a perpendicular, like @AkashdeepDeb explained.